# Best Agent Identification for General Game Playing

### v1

@@ Abstract
We present an efficient and generalised procedure to accurately identify the best performing algorithm for each sub-task in a multi-problem domain. Our approach treats this as a set of best arm identification problems for multi-armed bandits, where each bandit corresponds to a specific task and each arm corresponds to a specific algorithm or agent. We propose an optimistic selection process based on the Wilson score interval (Optimistic-WS) that ranks each arm across all bandits in terms of their potential regret reduction. We evaluate the performance of Optimistic-WS on two of the most popular general game domains, the General Video Game AI (GVGAI) framework and the Ludii general game playing system, with the goal of identifying the highest performing agent for each game within a limited number of trials. Compared to previous best arm identification algorithms for multi-armed bandits, our results demonstrate a substantial performance improvement in terms of average simple regret. This novel approach can be used to significantly improve the quality and accuracy of agent evaluation procedures for general game frameworks, as well as other multi-task domains with high algorithm runtimes.
@@
Keywords: General game playing, best arm identification, multi-armed bandits, Wilson score, agent evaluation
1 Introduction
@@
The field of general game playing focuses on the development of AI agents that are able to effectively play a wide variety of games (Genesereth et al., 2005). This can often be viewed as a multi-problem domain, where each agent needs to perform well across a collection of individual tasks. Designing artificial general agents that are able to successfully play a wide range of games has long been a core research area for artificial intelligence and machine learning, leading to significant breakthroughs in the areas of deep reinforcement learning (Torrado et al., 2018; Goldwaser and Thielscher, 2020), search and planning (de Waard et al., 2016; Frydenberg et al., 2015), transfer learning (Banerjee and Stone, 2007; Soemers et al., 2023), evolutionary algorithms (Gaina et al., 2017), among many others. However, as the number of games and agents developed for these frameworks increases, so too does the evaluation time needed to reliably assess the performance of each agent. Many agents do not play identically when repeatedly attempting a game, often requiring a large number of trials to obtain an accurate performance estimate (Bontrager et al., 2021). While there is often “no free lunch” when it comes to producing more results in less time (Ashlock et al., 2017), several approaches have been proposed to reduce the number of unnecessary evaluations using prior domain knowledge (Stephenson et al., 2020; Aitchison et al., 2023).
@@
Rather than needing to accurately assess the game playing abilities of each agent across all games, our aim is to develop a domain agnostic approach that can efficiently identify the best performing algorithm (i.e., agent) for each task (i.e., game) within a defined set. Knowing the best agent for each game can be useful for several purposes, such as training prediction models for portfolio/ensemble agents or detecting skill gaps in the existing agent suite (Anderson et al., 2019; Stephenson and Renz, 2017). We frame this challenge as a multi-bandit best arm identification problem (Gabillon et al., 2011), where each bandit represents a specific game and each arm represents an available agent for playing that game.
@@
In this paper, we propose a novel multi-bandit best arm identification algorithm called Optimistic-WS. This approach combines an “optimism in the face of uncertainty” philosophy with a generalised interpretation of the Wilson score interval, selecting arms that have the highest potential for improving our best arm estimation (i.e., minimising simple regret) across all bandits. We evaluate and compare the performance of our Optimistic-WS approach against prior multi-bandit best arm identification algorithms for two of the most popular general game domains. This includes the General Video Game AI (GVGAI) framework (Perez-Liebana et al., 2019) and the Ludii general game playing system (Piette et al., 2020), both of which are frequently used within the academic community for developing and comparing general game playing techniques (Perez-Liebana et al., 2016; Piette et al., 2023). Our results demonstrate that our proposed Optimistic-WS approach provides significant performance improvements with regards to best agent identification when compared to alternative state-of-the art techniques.
@@
The remainder of this paper is organised as follows: Section 2 formalises the multi-bandit best arm identification problem and its applicability to best agent identification for general game playing. Section 3 describes relevant prior work on general game systems, efficient agent evaluation procedures, best arm identification algorithms, and the Wilson score interval. Section 4 defines our proposed Optimistic-WS approach. Section 5 provides experiment details and performance evaluations using the GVGAI and Ludii general game playing domains. Section 6 summarises our results, highlighting potential areas for further investigation and improvement.
2 Problem Setup
@@
2.1 Multi-Bandit Best Arm Identification
Best-arm identification is a variation of the more traditional stochastic multi-armed bandit problem. Rather than maximising the cumulative reward over multiple arm pulls, the aim of best-arm identification is to maximise the reward of a single arm pull after a defined exploration period (i.e., select the arm with the highest expected reward after a limited number of prior arm pulls). This problem can be expanded further into the multi-bandit best arm identification problem, where instead of a single bandit machine there are now multiple bandits available to choose from, each with their own selection of arms and reward distributions. Rather than identifying the single best arm, the goal of multi-bandit best arm identification is to identify the best performing arm for each individual bandit. A more formalised definition of the multi-bandit best arm identification problem is provided below, with the majority of terms borrowed from (Gabillon et al., 2011).
@@ 2.1 Multi-Bandit Best Arm Identification
To evaluate the performance of our approach after $n$ rounds, for each bandit we compare the expected reward for our selected arm against that of the true best arm. This comparison is usually done by calculating the average difference in expected rewards for each arm pair, known as the simple regret $r(n)=\frac{1}{M}\sum_{m=1}^{M}(\mu_{m}^{\star}-\mu_{mJ_{m}(n)})$. In certain situations, the percentage of incorrectly identified best arms may be a more desirable measure of performance, $e(n)=\frac{1}{M}\sum_{m=1}^{M}\mathbb{P}(J_{m}(n)\neq k_{m}^{\star})$. For our case study, we will focus on using the average simple regret $r(n)$ as our measure of algorithm performance.
2.2 Best Agent Identification for General Game Playing
@@
The above problem of multi-bandit best arm identification is analogous to determining the best performing algorithms across a variety of different tasks. Each task can be treated as a single multi-armed bandit, with each arm representing a potential algorithm that can perform this task (albeit with varying degrees of success). Using this framing allows us to represent our task of identifying the best agent for each game within a general game framework as a multi-bandit best arm identification problem. Each game is treated as its own multi-arm bandit, with each arm of this bandit representing a different game playing agent. A pull of a specific arm for a specific bandit corresponds to running a single trial for the associated game-agent pair, with the agent’s outcome at the end of this trial determining the reward obtained.
2.2.1 Domain Considerations
While prior approaches for solving multi-bandit best arm identification problems have been proposed for the general case described above, our specific focus on general game playing introduces several additional considerations:
@@ 2.2.1 Domain Considerations
Reward Bounds - One assumption we can make for the domain of general game playing is that outcome for an agent at the end of a trial is always between 0 and 1 (inclusive). Typically an outcome of 0 indicates a loss and an outcome of 1 indicates a win, with other outcomes between these indicating either a draw or positional ranking among multiple players. This means that the value of $b$ in our reward distribution bounds will always be equal to 1 (i.e., all rewards bounded in $[0,1]$).
@@
Unequal Number of Arms - Another consideration is that the above definition of a multi-bandit environment assumes an equal number of arms $K$ for each bandit. However, in our application to general game playing it is often the case that certain agents may not be able to play every game available. In these situations, we can simply apply a generalisation to the above definitions that replaces the static value $K$ with a bandit-specific value $K_{m}$ that represents the number of arms available for a given bandit $m$.
@@
Multiple Best Arms - Lastly, when comparing multiple agents for a single game, the assumption that there will always be a unique best arm $k_{m}^{\star}$ for any bandit $m$ is unlikely to hold. For example, any game where several agents are always able to win will result in an equally high reward estimation of 1.0 for multiple arms. To account for this, we define our $\operatorname*{argmax}$ function to always return a single best argument combination, chosen at random from the highest value arguments. We will also need to modify the average probability of error measure $e(n)$, so that selecting any arm with the highest expected reward $\mu_{m}^{\star}$ will be recorded as a correct prediction (i.e., $e(n)=\frac{1}{M}\sum_{m=1}^{M}\mathbb{P}(\mu_{m}^{\star}\neq\mu_{mJ_{m}(n)})$).
3 Background
3.1 General Game Systems
@@
While many different general game domains have been proposed over the past decade, arguably the two most popular for academic research have been the General Video Game AI framework (GVGAI) and the Ludii general game playing system.
3.1.1 General Video Game AI Framework
@@
The General Video Game AI (GVGAI) framework contains over 100 simple arcade-style video games, described using the Video Game Description Language (VGDL) (Perez-Liebana et al., 2019). The GVGAI framework has been the subject of a long running general game playing competition (Perez-Liebana et al., 2016), leading to several dozen agents being developed for it over the years (Pérez-Liébana et al., 2016; Soemers et al., 2016; Gaina et al., 2017; Weinstein and Littman, 2012; Mendes et al., 2016). While there are several auxiliary research tracks available for the GVGAI framework, in this paper we will focus solely on the games and agents developed for the main single-player planning track. Each game includes five distinct levels, which all use the same base set of mechanics but vary the number and initial locations of objects. Previous analysis on agent performance across the corpus of GVGAI games has observed that agents often produce highly varied reward distributions (Bontrager et al., 2021). Many of the games available within the GVGAI framework also have some form of stochasticity, such as randomised items or enemy behaviour, making agent evaluation an inherently noisy process. These factors, along with the time consuming nature of running dozens of trials across thousands of game-agent pairs, makes the GVGAI framework an ideal case study for best agent identification.
3.1.2 Ludii general game playing system
@@
In contrast to the GVGAI framework’s focus on video games, the Ludii general game playing system instead provides over 1000 board and puzzle games (Piette et al., 2020). This makes Ludii one of the largest collections of playable games for AI research purposes. Games within Ludii are described using the ludemic game description language, which supports a wide range of mechanics and game types. This includes stochastic or hidden information, alternating or simultaneous moves, and player counts ranging from 1 to 16 (Stephenson et al., 2019). Ludii has also been used as a competition framework for promoting the development of general game playing agents (Piette et al., 2023), as well as research into general game heuristics (Stephenson et al., 2021). Much like the GVGAI framework, agent evaluation in Ludii is a lengthy and imperfect process, often requiring maximum turn limits to even guarantee that a game will end. These aspects make Ludii a prime candidate for applying more efficient best agent identification algorithms.
3.2 Efficient Agent Evaluation
@@ 3.2.1 Game Subsets
Rather than evaluating all agents across the full suite of available games, one alternative is to identify a representative subset of games to evaluate on. This approach has been previously investigated for both the General Video Game AI framework (Stephenson et al., 2020) and Arcade Learning Environment (Aitchison et al., 2023). This technique has also been mentioned as possible future work for the Ludii general game playing system (Piette et al., 2021; Stephenson et al., 2023) but has yet to be properly investigated. While this approach has a similar motivation to our current problem, being that testing all game-agent pairs a sufficient number of times to get a reliable indication of performance is computationally expensive, our goal is to instead focus on identifying the best agent for each individual game, rather than assessing the general performance of each agent. These approaches also use domain-specific measures of game and/or agent similarity to identify a representative game subset, whereas our presented approach requires no such domain knowledge.
3.2.2 Active Learning
@@
Active Learning is a branch of Machine Learning where labelled training instances are selected through an intelligent process (Settles, 2011). This typically occurs when there is a large quantity of unlabelled instances, but labelling any given instance is costly to perform. An active learning algorithm attempts to identify which instances should be labelled to provide maximal training benefit. This situation is highly similar to our described problem, but instead of an abundance of unlabelled instances that need labelling we instead have a noisy and/or probabilistic labelling process. In both cases, the core motivation that obtaining a large number of accurate instance labels is prohibitively expensive remains the same. While previous research has demonstrated that active learning approaches can be used to minimise reward estimate uncertainty for multi-armed bandits (Deng et al., 2011), the problem of best arm identification has not yet been explored.
3.3 Best Arm Identification
@@
Several prior algorithms for the problem of best arm identification have been presented over the past few decades, either focusing on the problem from a general perspective or for a specific set of circumstances. Common variations include correlated arm rewards (Gupta et al., 2021; Kazerouni and Wein, 2021; Hoffman et al., 2014), identifying the best-$N$ arms (Bubeck et al., 2013), identifying the best arm across multiple overlapping groups (Scarlett et al., 2019), and restless arms with evolving states (Narayana Prasad et al., 2022). The majority of the these approaches focus on best arm identification for a single bandit, but can also be generalised to a multi-bandit domain. Out of these past approaches, four algorithms (and variants of them) stand out as being the most applicable to our use case.
3.3.1 GapE

### v2

@@ Abstract
We present an efficient and generalised procedure to accurately identify the best (or near best) performing algorithm for each sub-task in a multi-problem domain. Our approach treats this as a set of best arm identification problems for multi-armed bandits, where each bandit corresponds to a specific task and each arm corresponds to a specific algorithm or agent. We propose an optimistic selection process based on a chosen confidence interval, that ranks each arm across all bandits in terms of their potential to influence our overall simple regret. We evaluate the performance of our approach on two of the most popular general game playing domains, the General Video Game AI (GVGAI) framework and the Ludii general game playing system, with the goal of selecting a high-performing agent for each game using a limited number of available trials. Compared to previous best arm identification algorithms for multi-armed bandits, our results demonstrate a substantial performance improvement in terms of average simple regret and average probability of error. This novel approach can be used to significantly improve the quality and accuracy of agent evaluation procedures for general game frameworks, as well as other multi-task domains with high algorithm runtimes.
Index Terms:
@@
General Game Playing, Best Arm Identification, Multi-Armed Bandits, Upper Confidence Bound, Agent Evaluation
[…]
=-15pt
I Introduction
@@
The field of general game playing focuses on the development of AI agents that are able to effectively play a wide variety of games [undeft]. This can often be viewed as a multi-problem domain, where each agent needs to perform well across a collection of individual tasks. Designing artificial general agents that are able to successfully play a wide range of games has long been a core research area for artificial intelligence and machine learning, leading to significant breakthroughs in the areas of deep reinforcement learning [undefav, undefu], search and planning [undefax, undefp], transfer learning [undefg, undefap], and evolutionary algorithms [undefr]. However, as the number of games and agents developed for these frameworks increases, so too does the evaluation time needed to reliably assess the performance of each agent. Many agents do not play identically when repeatedly attempting a game, often requiring a large number of trials to obtain an accurate performance estimate [undefi]. While there is often “no free lunch” when it comes to producing more results in less time [undefc], several approaches have been proposed to reduce the number of unnecessary evaluations using prior domain knowledge [undefar, undefa].
@@
Rather than needing to accurately assess the game playing abilities of each agent across all games, our aim is to develop a domain agnostic approach that can efficiently identify the best (or near best) performing algorithm (i.e., agent) for each task (i.e., game) within a defined set. Identifying a high-performing agent for each game can be useful for several purposes, such as training prediction models for portfolio/ensemble agents or detecting skill gaps in the existing agent suite [undefb, undefaq]. We frame this challenge as a multi-bandit best arm identification problem [undefq], where each bandit represents a specific game and each arm represents an available agent for playing that game. In this comparative setup, our goal is to select a single agent for each of our available games, with the goal of maximising the total reward (i.e., number of wins) across all games. To help us make our selection, we can run a limited number of exploratory trials to estimate the performance of each game-agent pair. The challenge therefore lies in determining which game-agent pairs (i.e., trials) should be run to help identify the top performing agent for each game.
@@
In this paper, we propose a novel multi-bandit best arm identification algorithm called RCP (Regret Change Potential). This approach combines an “optimism in the face of uncertainty” philosophy with a chosen confidence interval estimate, selecting arms that have the highest potential for minimising the overall simple regret across all bandits. We evaluate and compare the performance of our RCP approach with multiple confidence interval estimates, against several prior multi-bandit best arm identification algorithms for two of the most popular general game playing domains. This includes the General Video Game AI (GVGAI) framework [undefag] and the Ludii general game playing system [undefaj], both of which are frequently used within the academic community for developing and comparing general game playing techniques [undefaf, undefak]. Our results demonstrate that our proposed RCP approach provides significant performance improvements with regards to best agent identification, when compared to alternative state-of-the art techniques.
@@
The remainder of this paper is organised as follows: Section 2 formalises the multi-bandit best arm identification problem and its applicability to best agent identification for general game playing. Section 3 describes relevant prior work on general game systems, efficient agent evaluation procedures, best arm identification algorithms, and various confidence interval estimates. Section 4 defines our proposed RCP approach, and the motivation behind its design. Section 5 provides details on our experimental setup using the GVGAI and Ludii general game playing domains. Section 6 presents and discusses the relative performance of each best arm identification algorithm. Section 7 summarises our results, highlighting potential areas for further investigation and improvement.
II Problem Setup
@@
II-A Multi-Bandit Best Arm Identification
Best-arm identification is a variation of the more traditional stochastic multi-armed bandit problem. Rather than maximising the cumulative reward over multiple arm pulls, the aim of best-arm identification is to maximise the reward of a single arm pull after a defined exploration period (i.e., select the arm with the highest expected reward after a limited number of prior arm pulls). This problem can be expanded further into the multi-bandit best arm identification problem, where instead of a single bandit there are now multiple bandits available to choose from, each with their own selection of arms and reward distributions. Rather than identifying the single best arm, the goal of multi-bandit best arm identification is to identify the best performing arm for each individual bandit. Once the defined exploration period ends, a single arm pull of each bandit is performed and the rewards from each pull are summed to give the final result. The challenge of the multi-bandit setup comes from the need to effectively distribute the available exploratory pulls across all bandits, as well as the arms each bandit has, to obtain the necessary information to maximise the combined reward from a single pull of each bandit. A more formalised definition of the multi-bandit best arm identification problem is provided below, with the majority of terms borrowed directly from [undefq].
@@ II-A Multi-Bandit Best Arm Identification
To evaluate the performance of our approach after $n$ rounds, for each bandit we compare the expected reward for our selected arm against that of the true best arm. This comparison is usually done by calculating the average difference in expected rewards for each arm pair, known as the simple regret $r(n)=\frac{1}{M}\sum_{m=1}^{M}(\mu_{m}^{\star}-\mu_{mJ_{m}(n)})$. In certain situations, the percentage of incorrectly identified best arms may be a more desirable measure of performance, $e(n)=\frac{1}{M}\sum_{m=1}^{M}\mathbb{P}(J_{m}(n)\neq k_{m}^{\star})$. For our case study, we primarily focus on using the average simple regret $r(n)$ as our measure of algorithm performance, although additional performance comparisons using the average probability of error measure $e(n)$ are provided as well.
II-B Best Agent Identification for General Game Playing
@@
The above problem of multi-bandit best arm identification is analogous to determining the best performing algorithms across a variety of different tasks. Each task can be treated as a single multi-armed bandit, with each arm representing a potential algorithm that can perform this task (albeit with varying degrees of success). Using this framing allows us to represent our task of identifying the best agent for each game within a general game framework as a multi-bandit best arm identification problem. Each game is treated as its own multi-arm bandit, with each arm of this bandit representing a different game playing agent. A pull of a specific arm for a specific bandit corresponds to running a single trial for the associated game-agent pair, with the agent’s outcome at the end of this trial determining the reward obtained. While prior approaches for solving multi-bandit best arm identification problems have been proposed for the general case described above, our specific focus on general game playing introduces several additional domain considerations.
II-B1 Reward Bounds
@@
One assumption we can make for the domain of general game playing is that the outcome for an agent at the end of a trial is always between 0 and 1 (inclusive). Typically an outcome of 0 indicates a loss and an outcome of 1 indicates a win, with other outcomes between these indicating either a draw or positional ranking among multiple players. This means that the value of $b$ in our reward distribution bounds is always equal to 1 (i.e., all rewards bounded in $[0,1]$).
II-B2 Unequal Number of Arms
@@
Another consideration is that the above definition of a multi-bandit environment assumes an equal number of arms $K$ for each bandit. However, in our application to general game playing it is often the case that certain agents may not be able to play every game available. In these situations, we can simply apply a generalisation to the above definitions that replaces the static value $K$ with a bandit-specific value $K_{m}$ that represents the number of arms available for a given bandit $m$.
II-B3 Multiple Best Arms
@@
Lastly, when comparing multiple agents for a single game, the assumption that there is always a unique best arm $k_{m}^{\star}$ for any bandit $m$ is unlikely to hold. For example, any game where several agents are always able to win results in an equally high reward estimation of 1.0 for multiple arms. To account for this, we define our $\operatorname*{argmax}$ function to always return a single best argument combination, chosen at random from the highest value arguments. We also need to modify the average probability of error measure $e(n)$, so that selecting any arm with the highest expected reward $\mu_{m}^{\star}$ is recorded as a correct prediction (i.e., $e(n)=\frac{1}{M}\sum_{m=1}^{M}\mathbb{P}(\mu_{m}^{\star}\neq\mu_{mJ_{m}(n)})$).
III Background
III-A General Game Systems
@@
While many different general game playing domains have been proposed over the past decade, arguably the two most popular for academic research have been the General Video Game AI framework (GVGAI) and the Ludii general game playing system.
III-A1 General Video Game AI Framework
@@
The General Video Game AI (GVGAI) framework contains over 100 simple arcade-style video games, described using the Video Game Description Language (VGDL) [undefag]. The GVGAI framework has been the subject of a long running general game playing competition [undefaf], leading to several dozen agents being developed for it over the years [undefah, undefao, undefr, undefaz, undefad]. While there are several auxiliary research tracks available for the GVGAI framework, in this paper we focus solely on the games and agents developed for the main single-player planning track. Each game includes five distinct levels, which all use the same base set of mechanics but vary the number and initial locations of objects. Previous analysis on agent performance across the corpus of GVGAI games has observed that agents often produce highly varied reward distributions [undefi]. Many of the games available within the GVGAI framework also have some form of stochasticity, such as randomised items or enemy behaviour, making agent evaluation an inherently noisy process. These factors, along with the time consuming nature of running dozens of trials across thousands of game-agent pairs, makes the GVGAI framework an ideal case study for best agent identification.
III-A2 Ludii general game playing system
@@
In contrast to the GVGAI framework’s focus on video games, the Ludii general game playing system instead provides over 1000 board and puzzle games [undefaj]. This makes Ludii one of the largest collections of playable games for AI research purposes. Games within Ludii are described using the ludemic game description language, which supports a wide range of mechanics and game types. This includes stochastic or hidden information, alternating or simultaneous moves, and player counts ranging from 1 to 16 [undefas]. Ludii has also been used as a competition framework for promoting the development of general game playing agents [undefak], as well as research into general game heuristics [undefat]. Much like the GVGAI framework, agent evaluation in Ludii is a lengthy and imperfect process, often requiring maximum turn limits to even guarantee that a game ends. These aspects make Ludii a prime candidate for applying more efficient best agent identification algorithms.
III-B Efficient Agent Evaluation
@@ III-B1 Game Subsets
Rather than evaluating all agents across the full suite of available games, one alternative is to identify a representative subset of games to evaluate on. This approach has been previously investigated for both the General Video Game AI framework [undefar] and Arcade Learning Environment [undefa]. This technique has also been mentioned as possible future work for the Ludii general game playing system [undefai, undefau] but has yet to be properly investigated. While this approach has a similar motivation to our current problem, being that testing all game-agent pairs a sufficient number of times to get a reliable indication of performance is computationally expensive, our goal is to instead focus on identifying the best agent for each individual game, rather than assessing the general performance of each agent. These approaches also use domain-specific measures of game and/or agent similarity to identify a representative game subset, whereas our presented approach requires no such domain knowledge.
III-B2 Active Learning
@@
Active Learning is a branch of Machine Learning where labelled training instances are selected through an intelligent process [undefan]. This typically occurs when there is a large quantity of unlabelled instances, but labelling any given instance is costly to perform. An active learning algorithm attempts to identify which instances should be labelled to provide maximal training benefit. This situation is highly similar to our described problem, but instead of an abundance of unlabelled instances that need labelling we instead have a noisy and/or probabilistic labelling process. In both cases, the core motivation that obtaining a large number of accurate instance labels is prohibitively expensive remains the same. While previous research has demonstrated that active learning approaches can be used to minimise reward estimate uncertainty for multi-armed bandits [undefn], the problem of best arm identification has not yet been explored.
III-C Best Arm Identification
@@
Several prior algorithms for the problem of best arm identification have been presented over the past few decades, either focusing on the problem from a general perspective or for a specific set of circumstances. Common variations include correlated arm rewards [undefv, undefab, undefx], identifying the best-$N$ arms [undefk], identifying the best arm across multiple overlapping groups [undefam], and restless arms with evolving states [undefae]. The majority of the these approaches focus on best arm identification for a single bandit, but can also be generalised to a multi-bandit domain. Out of these past approaches, four algorithms (and variants of them) stand out as being the most applicable to our use case.
III-C1 GapE